OpenStudy (anonymous):

dx2/dx1 = x2(-3 + x1) / (x1 (-3 + 4x2) how do you solve by separating DE method?

OpenStudy (zehanz):

Assuming \(x_1\) and \(x_2\) are your variables, cross-multiplication is the first step: \(x_1(-3+4x_2)dx_2=x_2(-3+x_1)dx_1\). Now divide by \(x_1\) and \(x_2\): \(\dfrac{-3+4x_2}{x_2}dx_2=\dfrac{-3+x_1}{x_1}dx_1\) Variables are separated now.

OpenStudy (zehanz):

To solve, you'd better split the fractions like this:\[\int\limits \left(\frac{-3}{x_2}+4\right)dx_2=\int\limits \left(\frac{-3}{x_1}+1\right)dx_1+C\]This will be equal to:\[-3 \ln|x_2|+4x_2=-3 \ln|x_1|+x_1+C\] To get \(x_2\) as a function of \(x_1\) from this formula doesn't look easy though...

OpenStudy (anonymous):

Thank you!!! i get it now!!

OpenStudy (zehanz):

YW!